# OC3D: Weakly Supervised Outdoor 3D Object Detection with Only Coarse Click Annotation

### v2

@@ Introduction
Referencing (Yang, Fan, and Zhang 2024), we redesigned the supervision assignment strategy. For all box-level pseudo-labels $\mathbf{A}_{b}$ of size $A_{b}$, we train the detection network to focus on its position, shape, orientation, and semantics. For mask-level pseudo-labels $\mathbf{A}_{m}$ of size $A_{m}$, the detector focuses solely on semantics and the center of masks. Therefore, the loss function for OC3D can be formulated as:
$\mathfrak{L}=\frac{1}{A_{b}}\sum_{A_{b}}\mathfrak{L}_{\operatorname{reg}}+\frac{1}{A_{b}+A_{m}}\sum_{A_{b}+A_{m}}\mathfrak{L}_{\operatorname{cls}}+\lambda\frac{1}{A_{m}}\sum_{A_{m}}\mathfrak{L}_{\operatorname{pos}}.$ (5)
@@
$\mathfrak{L}_{\operatorname{reg}}$ and $\mathfrak{L}_{\operatorname{cls}}$ are commonly used regression and classification losses in 3D object detection. $\mathfrak{L}_{\operatorname{pos}}$ is the part that decouples the central position from $\mathfrak{L}_{\operatorname{reg}}$. Since mask centers are not accurate instance centers, we set the hyper-parameter $\lambda$ to reduce the weight of this part of the loss.
Mask2Box Enhanced Training.
Mask-level pseudo-labels provide only semantic and coarse localization information, lacking a description of the instance shape. To compensate for the information lost in mask-level pseudo-labels, we use high-confidence bounding-box predictions to upgrade mask-level pseudo-labels. As shown in Fig. 3, based on mask-level pseudo-labels, we use the box fitting algorithm (Zhang et al. 2017) to generate a temporary bounding box (Temp. Box). Then, we calculate the IoU of the temporary box with all high-confidence predictions and return the prediction box with the highest IoU:
$\operatorname{Proposal}=\underset{\operatorname{IoU}}{\operatorname{arg\,max}}\,(\operatorname{IoU}_{i}=\frac{\operatorname{Temp.Box}\cap l_{i}}{\operatorname{Temp.Box}\cup l_{i}}).$ (6)
where $l_{i}$ originates from all the high-confidence predictions. If the point-level pseudo-label is within the proposal, then the proposal is used to replace the point-level pseudo-label. Conversely, the point-level pseudo-label is retained and will be awaited for the next upgrade opportunity.
Figure 3: An illustration of the upgrading strategy for mask-level pseudo labels.
Method Venue Annotations 3D-Detection BEV-Detection
Type Cost Easy Mod Hard Easy Mod Hard
Voxel-RCNN (Deng et al. 2021) AAAI2021 Fully Supervised 100% 98.7 94.9 94.5 98.7 94.9 94.6
WS3D (Meng et al. 2021) PAMI2021 Boxes + Center Clicks 3% 96.3 89.4 88.9 96.4 89.3 88.9
MixSup (Yang et al. 2024) ICLR2024 94.9 92.7 90.0 94.9 93.0 90.4
SS3D (Liu et al. 2022a) CVPR2022 Sparse Boxes 2% 98.3 89.2 88.3 - - -
CoIn (Xia et al. 2023b) ICCV2023 96.3 86.7 74.4 96.3 88.4 75.6
CoIn++ (Xia et al. 2023b) ICCV2023 99.3 92.7 88.8 - - -
HINTED (Xia et al. 2024) CVPR2024 98.5 91.6 90.3 98.4 92.9 90.6
OC3D++ - Sparse Coarse Clicks 0.2% 96.4 91.6 84.6 96.5 92.0 84.9
OC3D - Coarse Clicks 1% 96.6 92.5 91.5 96.7 94.1 92.0
Table 1: Experimental results on KITTI dataset compared with recent state-of-the-art label-efficient methods. We report results of ’car’ with 40 recall positions, below the 0.5 IoU thresholds. ’Boxes + Center-Clicks’ denotes that in some scenarios, bounding box annotations are retained, while in the rest, only the central position annotations are preserved. ’Sparse Boxes’ indicates that only one bounding box annotation is retained per ten scenes.
@@
Method Annotation mAP NDS Car Truck C.V. Bus Trailer Barrier Motor. Bike Ped. T.C.
CenterPoint Fully Supervised (100%) 56.18 64.69 84.10 54.56 16.38 67.31 36.95 65.27 53.58 35.76 82.70 65.08
CenterPoint Sparse Boxes (2%) 8.09 25.77 24.62 2.84 0.00 15.66 0.00 4.07 3.33 0.29 25.11 4.96
CoIn 12.47 33.79 38.70 6.85 0.00 20.67 7.81 11.51 2.85 3.36 34.85 8.50
HINTED 32.62 45.76 66.63 32.71 7.59 54.56 11.42 21.16 29.12 19.08 57.22 26.63
OC3D++ Sparse Coarse Clicks (0.2%) 24.88 38.12 56.37 23.47 1.10 33.30 3.43 16.18 24.67 8.81 54.42 27.00
OC3D Coarse Clicks (1%) 44.04 49.87 78.91 46.18 8.01 50.34 27.68 60.70 28.80 10.70 74.06 54.98
Table 2: The multi-class results on the nuScenes val set. ‘C.V.’, ‘Ped.’, and ‘T.C.’ are short for construction vehicle, pedestrian, and traffic cones, respectively.
Exploration of Sparse Click 3D Detector
Recently, weakly supervised 3D object detection methods (Xia et al. 2023b, 2024) with sparse bounding boxes have made great progress. The sparse annotation approach can greatly reduce the annotation cost. Motivated by this, we further explore the weakly supervised 3D object detection scheme based on sparse clicks. Unlike conventional click annotations, under sparse clicking, only one instance is randomly selected for annotation in each scene, presenting a greater challenge to the 3D object detector. To address the problem of missing annotation information, we propose an OC3D-based missing-click instances mining method, OC3D++ (for specific design details, see the supplementary materials). The experimental results of OC3D++ demonstrate that our OC3D can be effectively combined with missing-click instances mining methods.
Experiments
@@
For KITTI dataset (Geiger, Lenz, and Urtasun 2012), to acquire a dense point cloud, we extracted point cloud data from consecutive frames of the raw data. It is worth noting that the raw data is only used to assist in the generation of pseudo-labels and does not participate in the training process of the detector. For the training stage, following traditional fully-supervised methods (Shi et al. 2020; Zhou and Tuzel 2018; Xia et al. 2023a), KITTI 3D detection dataset is divided into a $train$ split of $3712$ frames and a $val$ split of $3769$ frames. For the $train$ split, we only retain the coarse click annotations. Moreover, we further generate an extremely low-cost training set (denoted as the $limited$ split), retaining the click annotations for only one instance in each training scene. To ensure a fair comparison, following sparsely-supervised methods (Liu et al. 2022a; Xia et al. 2023b), we calculate the mean Average Precision (mAP) using 40 recall positions.
NuScenes (Caesar et al. 2020) is a comprehensive autonomous driving dataset that includes 1000 sequences, consisting of 700, 150, and 150 sequences for training, validation, and testing, respectively. There are a total of 28k annotated frames for training, 6k for validation, and another 6k for testing. We also retain coarse click annotations for all objects in the $train$ split in nuScenes. Considering multi-class sample balance, we performed a single click annotation for each category present in the scene to obtain $limited$ split for nuScenes. For evaluation, we adopt mAP and nuScenes detection score (NDS) as the main metrics.
@@
For coarse click annotations, we use the BEV center of ground truth as a reference and apply a large perturbation range ($0.5\times w,0.5\times l$) to simulate coarse manual clicks. In the supplementary material, we demonstrate a comparison between the effects of simulated clicks and real coarse manual clicks. Meanwhile, in Tab.4, we analyze the impact of different perturbation ranges on the results. Due to the large number of categories in nuScenes and the significant size differences between them, we set different neighborhood radii $r$ for different categories based on common human prior knowledge. Specific settings can be found in the supplementary material.
@@
We are the first to develop a method for training object detectors only with click-annotations, and there are no previously published baselines for comparison. To validate the effectiveness of our OC3D, we chose to compare it with works that are also label-efficient. To ensure a fair comparison, we adopt widely used VoxelRCNN (Deng et al. 2021) as basic detector architectural. Furthermore, leveraging the proposed Click2Box module, we infer bounding box supervision from click annotations to construct multiple baselines with click annotations.
Comparison with State-of-the-art
Validation on KITTI.
In Tab. 1, we conducted experiments to compare our approach with state-of-the-art label-efficient methods on KITTI. Following the mainstream approaches (Liu et al. 2022a; Xia et al. 2023b, 2024), we also adopted Voxel-RCNN (Deng et al. 2021) as the base detector. For coarse click annotations, the labeling time per instance is approximately 1.2 seconds, which is about 2 times faster than center-click labeling and 100 times faster than bounding box labeling. Despite employing a more lightweight annotation form, retaining only coarse click annotation, our OC3D still achieves comparable performance with other methods. In the Car-3D detection task, at the ‘Easy’ and ‘Mod’ difficulty levels, the gap between our performance and the previous best is only $3.3\%$ and $0.2\%$, respectively. Meanwhile, at the ’Hard’ difficulty level, our method achieved the best performance. In the Car-BEV detection task, at the ’Easy’ difficulty level, the gap between our performance and the previous best is only $2.3\%$. Meanwhile, at the ’Hard’ difficulty level, our method achieved the best performance. Especially, under conditions of an extremely cost-effective sparse clicking, the proposed OC3D++ demonstrates the capability to sustain consistent performance.
Validation on NuScenes.
We also conducted validation experiments on the nuScenes dataset. To ensure a fair comparison, we follow the previous methods (Xia et al. 2023b, 2024) to select the CenterPoint (Yin, Zhou, and Krähenbühl 2021) as the base detector. In Tab. 2, where OC3D achieves performance comparable to the fully supervised baseline method. This confirms that our SC3D still demonstrates outstanding performance on highly challenging multi-class tasks. In the experimental with sparse coarse click annotations, the mAP of OC3D++ decreased. The reason is the higher number of instances in the nuScenes, which poses a greater challenge for sparse settings.
Ablation Study and Analysis
MST MET Car AP@IoU 0.5 Car AP@IoU 0.7
@@
86.3 88.6 43.6 70.3
87.7 90.0 51.0 71.3
92.5 94.1 66.3 82.9
Table 3: Ablation study under different module. We report on the Car-mod results with 40 recall positions, below the 0.5 and 0.7 IoU thresholds.
$\delta$ Car-3D AP@IoU 0.5 Car-BEV AP@IoU 0.5
@@
Figure 4: Prediction results and the quality of pseudo-labels across various iterative rounds. (a) and (b) represent the 3D and BEV detection results for cars at IOU thresholds of 0.5 and 0.7, respectively; (c) compare the box-level pseudo-labels with the ground truth across each iterative round.
Effect of Mixed Supervised Training (MST) and Mask2Box Enhanced Training (MET).
The effects of the proposed MST, and MET are listed in Tab. 3. The first row of Tab. 3 presents the baseline performance, using only the proposed Click2Box strategy to generate bounding box pseudo-labels from click annotations. The results in the second row represent the detector performance by mixed supervised training with mixed pseudo-labels. Our design eliminates the adverse effects of dynamic low-quality bounding box pseudo-labels and achieves better detection results. Finally, based on Click2Box and Click2Mask, we designed the Mask2Box module to continuously upgrade mask-level pseudo-labels to box-level pseudo-labels, providing more accurate supervision signals that significantly improve detection accuracy.
Effects of Perturbation Factor $\delta$ and Weight Selection for $\lambda$.
Tab. 4 demonstrates the robustness of the coarse click. To rigorously verify the impact of the perturbation factor, we conduct experiments on the more challenging sparse clicking setup. We applied various disturbance factors $\delta=0.25,0.5,0.7,1.0$, defining the disturbance range as $(\delta\times w,\delta\times l)$ , where $w$ and $l$ are the size of the object. Even with a large disturbance range ($\delta=1.0$), our OC3D++ still achieves satisfactory performance. In Tab.5, we discuss the hyper-parameter $\lambda$ in Equ.(5) without the Mask2Box modules. Lower weights for ambiguous location predictions yield better results. This result indicates that for ambiguous location supervision, providing a single hint is sufficient to achieve optimal detection performance, and excessive weighting introduces noise to the detector.
$\lambda$ Car-3D AP@IoU 0.5 Car-3D AP@IoU 0.7
@@
Table 5: Weights Selection for $\lambda$. We report results on the Car-3D detection with 40 recall positions, below the 0.5 and 0.7 IoU thresholds.
@@
We show the performance of OC3D++ over different numbers of training iterations in Fig. 4 (a) and (b). At both the 0.5 and 0.7 thresholds, the performance is significantly enhanced. Using the ground truth as a reference, we observed the average IoU and recall throughout the iterative learning process. As shown in Fig. 4 (c), there is an overall stable upward trend in the quality of our pseudo-labels as the number of iterations increases. Meanwhile, we also observed a decrease in IoU and a sudden increase in Recall during the 5th iteration. This trend is attributed to the retention of additional pseudo-labels, which is a normal occurrence in the iterative process.
Discussion and Conclusion
We designed an efficient annotation strategy called coarse click annotation for 3D object detection and proposed a weakly-supervised object detection method, OC3D, leveraging this approach. In the mixed pseudo-label generation stage, we propose a novel method for dynamic and static classification, and design the Click2Box and Click2Mask modules according to the motion states of objects to generate mixed pseudo-labels. Subsequently, in the Mask2Box enhanced mixed supervised training stage, We train the 3D detector with mixed pseudo-labels and design the Mask2Box module to obtain richer supervisory information. Extensive experiments on KITTI and nuScenes have shown that our OC3D achieves commendable performance with purely clicks. Even in the more challenging sparse clicking, our OC3D++ continues to deliver outstanding performance.

### v3

@@ 3.2 Mixed-Supervised Teacher Training
To optimize the initial teacher detector, we redesigned the supervision assignment strategy. For all box-level pseudo-labels $\mathbf{L}_{b}$, we train the detection network to focus on its position, shape, orientation, and semantics. For mask-level pseudo-labels $\mathbf{L}_{m}$, the detector focuses solely on semantics and the center of masks. Therefore, the loss function for teacher network can be formulated as:
$\mathfrak{L}_{mix}=\frac{1}{|\mathbf{L}_{b}|}\sum_{\mathbf{L}_{b}}\mathfrak{L}_{\operatorname{reg}}+\frac{1}{|\mathbf{L}_{b+m}|}\sum_{\mathbf{L}_{b+m}}\mathfrak{L}_{\operatorname{cls}}+\frac{\lambda}{|\mathbf{L}_{m}|}\sum_{\mathbf{L}_{m}}\mathfrak{L}_{\operatorname{pos}}.$ (5)
@@
$\mathfrak{L}_{\operatorname{reg}}$ and $\mathfrak{L}_{\operatorname{cls}}$ are commonly used regression and classification losses in 3D object detection. $\mathfrak{L}_{\operatorname{pos}}$ is the part that decouples the central position from $\mathfrak{L}_{\operatorname{reg}}$. $\mathbf{L}_{b+m}$ is the union of ${\color[rgb]{1,.5,0}\mathbf{L}_{b}}$ and ${\color[rgb]{0.546875,0,0}\mathbf{L}_{m}}$. Since mask centers are not accurate instance centers, we set the hyper-parameter $\lambda$ to reduce the weight of this part.
Mask2Box Enhanced Teacher.
Mask-level pseudo-labels provide only semantic and coarse localization information, lacking a description of the instance shape. To compensate for the information lost in mask-level pseudo-labels, we use high-confidence bounding-box predictions of initial teacher to upgrade mask-level pseudo-labels. Specifically, we first set a threshold $\tau$ to filter out low-confidence predictions from the detector’s output. Then, we establish a matching relationship based on the intersection-over-union (IoU) between the ${\color[rgb]{0.546875,0,0}\mathbf{L}_{m}}$ and high-quality prediction boxes $Y$. The bounding box with the highest IoU would replace the ${\color[rgb]{0.546875,0,0}\mathbf{L}_{m}}$ to participate in subsequent training. Mask2Box progressively refines mixed supervision, further enhancing the performance of the teacher network.
3.3 Mixed-Supervised Student Training
Figure 4: Training strategies for the mixed-supervised student network.
By progressively recovering from click annotations to supervision signals, we enhance the performance of the teacher model. However, the abundance of unclicked instances within the scenes impedes the training of a well-performed detector. Inspired by  [43], we utilize the inherent generalization ability of the teacher network to mine unlabeled instances. Following this idea, ensuring the accuracy of the pseudo-labels output by the teacher model is our primary concern.
Transformation Equivariant Guided Evaluation.
The pseudo-label evaluation solely based on prediction confidence cannot guarantee the reliability of predicted labels [30], which also limits the performance of the teacher network. Inspired by [38, 16, 33], we utilize the transformation equivariant of the detection network to evaluate predicted labels. Specifically, as shown in Fig. 4, we first perform global augmentation (including random rotation, flipping, and scaling) on the input point cloud, and the teacher model produces corresponding detection results $\hat{Y}$. Then, we inversely transform the augmented results and align them back to the initial coordinate system. Based on the property of transformation equivariant, we consider the bounding boxes with high alignment between the $\hat{Y}$ and the original high-quality predicts ${Y}$ as promising pseudo-labels. We use the pairwise IoU scores between $\hat{Y}$ and ${Y}$ as the pseudo-label alignment scores.
Mixed-supervision Expanding.
To further refine label quality, we adopt a dual-threshold filter strategy [18], which uses cluster operation to generate two thresholds $\mu_{h}$ and $\mu_{l}$ based on the distribution of IoU scores. For pseudo-labels with alignment scores greater than the high threshold $\mu_{h}$, we directly utilize them to provide supervision for the training of the student network and serve as support for the Mask2Box module. For pseudo-labels with alignment scores between the two thresholds, they can participate in the training of the student network as mask-level supervision. For pseudo-labels with alignment scores below the low threshold $\mu_{l}$, we will discard them directly. The training loss of the student network is consistent with that of the teacher network.
4 Experiments
@@
Method Annotation (cost) mAP NDS Car Truck C.V. Bus Trailer Barrier Motor. Bike Ped. T.C.
CenterPoint [47] Fully Supervised (100%) 56.18 64.69 84.10 54.56 16.38 67.31 36.95 65.27 53.58 35.76 82.70 65.08
WS3D [24] Boxes(10%) + Center Clicks 39.52 47.81 78.59 42.79 5.15 38.60 17.96 46.38 27.45 16.98 75.89 45.37
MixSup [46] Boxes(10%) + SAM [14] 49.49 58.65 64.63 41.71 15.61 57.57 28.19 43.56 62.28 51.42 75.07 54.87
CenterPoint [47] Sparse Boxes (2%) 8.09 25.77 24.62 2.84 0.00 15.66 0.00 4.07 3.33 0.29 25.11 4.96
CoIn [42] 12.47 33.79 38.70 6.85 0.00 20.67 7.81 11.51 2.85 3.36 34.85 8.50
HINTED [43] 32.62 45.76 66.63 32.71 7.59 54.56 11.42 21.16 29.12 19.08 57.22 26.63
SC3D-Teacher Sparse Coarse Clicks (0.2%) 24.88 38.12 56.37 23.47 1.10 33.30 3.43 16.18 24.67 8.81 54.42 27.00
SC3D-Student 30.71 42.37 64.46 30.73 5.21 45.21 10.37 19.26 27.33 8.35 59.64 36.59
Table 1: The multi-class results on the nuScenes val set. ‘C.V.’, ‘Ped.’, and ‘T.C.’ are short for construction vehicle, pedestrian, and traffic cones, respectively. ’Boxes + Center-Clicks’ denotes that in some scenarios, bounding box annotations are retained, while in the rest, only the central position annotations are preserved. ’Sparse Boxes’ indicates that only one bounding box annotation is retained per ten scenes.
@@
NuScenes [1] is a comprehensive autonomous driving dataset that includes 1000 sequences, consisting of 700, 150, and 150 sequences for training, validation, and testing, respectively. There are a total of 28k annotated frames for training, 6k for validation, and another 6k for testing. We perform sparse click annotation on the $train$ split of nuScenes to get a new $train$ split. For evaluation, we adopt mAP and nuScenes detection score (NDS) as the main metrics.
For KITTI dataset [10], to acquire a dense point cloud, we extracted point cloud data from consecutive frames of the raw data. It is worth noting that the raw data is only used to assist in the generation of pseudo-labels and does not participate in the training process of the detector. For the training stage, following traditional fully-supervised methods [28, 53, 41], KITTI 3D detection dataset is divided into a $train$ split of $3712$ frames and a $val$ split of $3769$ frames. We perform sparse click annotation on the $train$ split of KITTI to get a new $train$ split. To ensure a fair comparison, following sparsely-supervised methods [16, 42], we calculate the mean Average Precision (mAP) using 40 recall positions with IoU@0.5.
@@
For coarse click annotations, we use the BEV center of ground truth as a reference and apply a large perturbation range ($0.5\times w,0.5\times l$) to simulate coarse manual clicks. In the supplementary material, we demonstrate a comparison between the effects of simulated clicks and real coarse manual clicks. Meanwhile, in Tab.6, we analyze the impact of different perturbation ranges on the results. Drawing on human prior knowledge, we establish multi-scale search radii to accommodate objects of different categories. At the training stage, we trained SC3D with a batch size of 32 and a learning rate of 0.003 for eighty epochs on 4 RTX 3090 GPUs. More details will be provided in the appendix.
@@
We are the first to develop a method for training object detectors only with click-annotations, and there are no previously published baselines for comparison. To validate the effectiveness of our SC3D, we chose to compare it with works that are also label-efficient. To ensure a fair comparison, we adopt widely used VoxelRCNN [5] as basic detector architectural. Furthermore, leveraging the proposed Click2Box module, we infer bounding box supervision from click annotations to construct multiple baselines with click annotations.
4.1 Main Results
Comparison with State-of-the-art Methods.
For coarse click annotations, the labeling time per instance is approximately $1.2$ seconds, which is about $2$ times faster than center-click labeling and 100 times faster than bounding box labeling. We primarily conduct comparative experiments with state-of-the-art methods on the nuScenes and KITTI datasets.
At first, we conducted experiments to compare our approach with state-of-the-art label-efficient methods on the nuScenes dataset. To ensure a fair comparison, we follow the previous methods [42, 46] to select the CenterPoint [47] as the base detector. SC3D-Teacher and SC3D-Student, respectively, represent our mixed-supervised teacher detector and mixed-supervised student detector. As shown in Tab. 1, SC3D-student achieves performance comparable to the previous best sparsely-supervised method. Meanwhile, the annotation cost of our SC3D-student reduces by more than $10$ times compared to the previously best-performing method. Compared to the precision of the teacher detector, the mAP of the student detector has increased by $5.83\%$, indicating that we have successfully leveraged the teacher detector to mine a large amount of unlabeled instance information. Due to our bounding box pseudo-labels being fitted based on clustering results, our approach performs poorly on larger-scale categories in the nuScenes dataset.
Method Annotations (cost) 3D-Detection
Easy Mod Hard
Voxel-RCNN [5] Fully Supervised 100% 98.7 94.9 94.5
WS3D [24] Boxes + Center Clicks (3%) 96.3 89.4 88.9
MixSup [46] 94.9 92.7 90.0
SS3D [16] Sparse Boxes (2%) 98.3 89.2 88.3
CoIn [42] 96.3 86.7 74.4
CoIn++ [42] 99.3 92.7 88.8
HINTED [43] 98.5 91.6 90.3
SC3D-Teacher Sparse Coarse Clicks (0.2%) 94.7 87.4 80.3
SC3D-Student 96.4 91.6 84.6
Table 2: Experimental results on KITTI dataset compared with recent state-of-the-art label-efficient methods. We report results of car with 40 recall positions, below the 0.5 IoU thresholds.
For KITTI dataset, following the mainstream approaches [16, 42, 43], we also adopted Voxel-RCNN [5] as the base detector. Since our initial bounding box labels are generated based on point cloud distribution and significantly differ from manual annotations, we refer to [48] and compare the evaluation results under the $0.5$ IOU threshold. Despite employing a more lightweight annotation form, retaining only coarse click annotation, our SC3D still achieves comparable performance with other methods. Furthermore, with only $0.2\%$ of the labeling cost, the proposed SC3D achieves $94\%$ of the average performance of full supervision.
Comparison with the mixed-annotation method.
Method Annotation 3D AP@0.7 BEV AP@0.7
WS3D clicks + $precisely^{\#}$ 84.0 75.1 73.2 88.5 84.9 84.7
SC3D s-clicks 69.3 56.0 50.0 89.1 76.9 69.6
SC3D s-clicks + $precisely^{\#}$ 91.6 77.4 74.3 95.4 86.1 83.3
Table 3: Comparison with WS3D on KITTI $val$ split. We report results of car with 40 recall positions, below the 0.7 IoU thresholds. “s-clicks” denotes sparse coarse clicks. $precisely^{\#}$ denotes 534 precisely-annotated instances.
To reduce the impact of manual annotation habits on label quality, we refer to mixed annotation strategies[24], incorporating a small number of additional manually annotated bounding boxes during detector training. As shown in Tab. 3, with the same number of precisely annotated instances, our approach can achieve better performance than WS3D at an IoU threshold of 0.7 using fewer clicks. This proves that SC3D can learn human annotation habits with a small amount of manual annotation, ultimately reaching the goal of adjusting bounding boxes.
Comparison with Full Clicks.
To validate the detection performance of SC3D on unlabeled instances, we introduced a new click setting for performance comparison: Full Coarse Clicks, which involves annotating clicks for all instances in the nuScenes and KITTI $train$ splits. Tab. 4 displays the experimental results conducted with two distinct click annotation schemes. From the experimental results, it is evident that compared to full click annotations, SC3D achieves similar performance under sparse annotations, especially on the KITTI dataset, with a precision loss of no more than $2\%$.
Method Annotation nuScenes KITTI
mAP NDS 3D BEV
SC3D Full Coarse Clicks 44.0 49.8 92.5 94.1
Sparse Coarse Clicks 30.7 42.3 91.6 92.0
Table 4: Comparison with the results of different click cost. Full Coarse Clicks refers to the rough annotation of all instances in the scene through coarse clicking. We report the results of mAP and NDS for the nuScenes dataset, as well as 3D@0.5 and BEV@0.5 for the KITTI dataset.
Figure 5: Prediction results and the quality of pseudo-labels across various iterative rounds. (a) and (b) represent the 3D and BEV results for cars at IOU thresholds of 0.5 and 0.7, respectively; (c) compare the pseudo-labels with the ground truth across each iterative round.
4.2 Ablation Study and Analysis
In this section, we conduct ablation studies to evaluate our proposed modules and parameters in SC3D. Models are trained on KITTI’s training set and tested on its validation set. We use Voxel-RCNN for our study due to its quick training, and results with other detectors are comparable.
MPLG Mask2Box MSS Car AP@0.5 Car AP@0.7
@@ 4.2 Ablation Study and Analysis
69.5 79.3 17.6 49.8
87.4 90.0 26.1 53.0
89.4 91.6 53.0 73.9
91.6 92.0 56.0 76.9
Table 5: Ablation study under different module. We report the car-mod. results with 40 recall positions, under 0.5 and 0.7 IoU thresholds.
Effect of Mixed Pseudo Label Generation (MPLG).
Tab. 5 presents the results of testing each module individually. The first row shows the results of generating box-level pseudo-labels directly from click annotations using the unsupervised approach [51]. Due to the difficulty in recovering the geometric structure of moving objects across consecutive frames, naive label generation methods struggle to ensure label quality. The second row demonstrates the outcomes of our proposed MPLG module, significantly enhancing the performance of the baseline method. This confirms the effectiveness of MPLG’s strategy for generating corresponding labels based on the motion status of objects.
Effect of Mask2Box.
Although the mask-level labels generated by MPLG can convey supervision information about moving objects, this descriptive form still cannot fully express the structure of moving objects. To address this issue, we designed the Mask2Box module to enhance mask-level pseudo-labels. The results in the third row of Tab. 5 validate the effectiveness of Mask2Box, particularly with a substantial increase in accuracy at an IoU threshold of $0.7$.
Effect of Mixed-Supervised Student (MSS) training.
The teacher-student network has been proven to be a viable solution for addressing sparsity issues [43]. Therefore, we propose a mixed-supervised teacher-student network for sparse click strategies, where the teacher detector leverages transformation equivariance to provide high-quality labels for the student network. The last row of Tab. 5 demonstrates the results of the student network, indicating that the teacher-student network architecture has further enhanced the performance of the detector.
Effects of Perturbation Factor $\delta$.
$\delta$ Car-3D AP@ 0.5 Car-BEV AP@ 0.5
@@
Tab. 6 demonstrates the robustness of the coarse click. To rigorously verify the impact of the perturbation factor, we conduct experiments on the more challenging sparse clicking setup. We applied various disturbance factors $\delta=0.25,0.5,0.7,1.0$, defining the disturbance range as $(\delta\times w,\delta\times l)$ , where $w$ and $l$ are the size of the object. Although the performance of the detector experiences a slight decline with the increase of perturbation, the overall performance remains relatively stable. Even with a large disturbance range ($\delta=1.0$), our SC3D still achieves satisfactory performance.
Weight Selection for $\lambda$.
In Tab. 7, we delve into the discussion of the hyper-parameter $\lambda$ as presented in Equation (5), specifically in the context of our model without the incorporation of the Mask2Box modules. The results of Tab. 7 reveal that assigning lower weights to predictions with ambiguous locations leads to improved outcomes. This observation suggests that, in the case of ambiguous location supervision, supplying a single, subtle cue is adequate to attain peak detection performance. Conversely, assigning excessive weight to such ambiguous cues can introduce unwanted noise into the detection process, potentially degrading the model’s accuracy and reliability.
$\lambda$ Car-3D AP@ 0.5 Car-3D AP@ 0.7
@@
Table 7: Weights Selection for $\lambda$. We report results with 40 recall positions, below the 0.5 and 0.7 IoU thresholds.
@@
In Fig. 5 (a) and (b), we show the performance of SC3D over different numbers of updates and iterations in the teacher-student network. At both the 0.5 and 0.7 thresholds, the performance is significantly enhanced. The improvement may be attributed to the student network receiving more information from unlabeled instances during the iterative updates. To verify this hypothesis, we use the ground truth as a reference, obseving the average IoU and recall throughout the iterative learning process. As shown in Fig. 5 (c), there is an overall stable upward trend in the quality of our pseudo-labels as the number of iterations increases. The increase in Recall confirms the successful mining of unlabeled instances.
5 Discussion and Conclusion
We designed an efficient annotation strategy, single click per frame, tailored for label-efficient 3D object detection and proposed a weakly-supervised object detection method, SC3D, leveraging this approach. SC3D is mainly divided into three parts: mixed label generation, mixed teacher model training, and mixed student model training. In the mixed pseudo-label generation stage, we propose a novel method for dynamic and static classification, and design the Click2Box and Click2Mask modules according to the motion states of objects to generate mixed pseudo-labels. Subsequently, we propose a mixed-supervised teacher model that empowers the detector to assimilate a blend of supervision signals. Ultimately, we introduce a mixed-supervised student network, which leverages the teacher model’s generalized insights to learn from unclicked instances. Extensive experiments on the nuScenes and KITTI datasets have demonstrated that our SC3D method, using only sparse clicks and incurring a labeling cost of just 0.2%, can achieve commendable performance.